# Epidemic-induced local awareness behavior inferred from surveys and genetic sequence data

### v2

@@ Discussion
As with any trade-off, our proposed analysis comes with a number of limitations. We identify superspreading events based on simple thresholding of sequence counts, which is less accurate than manual contact tracing, where more metadata and more context about infection events can be taken into account. Consequently, we only compute highly aggregated statistics on the detected events. One ECS gives only very noisy information about the outcome of each superspreading event, and only the median of all ECSs, the $\mathrm{ECS}_{c}$ value has the statistical power to say anything about local awareness in region $c$. Since the amount of genetic sequences we have available since COVID-19 is unprecedented, and the new tools to analyze it are just being developed [23], our results too have to be confirmed by further research.
@@
From an operational perspective, by studying $\mathrm{ECS}_{c}$ values, we are able to measure how effectively different countries managed to contain superspreading events in different waves. We observe that this effectiveness is highly correlated with the containment policies implemented in each country, suggesting that stricter government policies could motivate the public to undertake stricter voluntary prevention methods. We envision that similar analyses will be used to evaluate the effectiveness of the implemented policies in future pandemics, potentially generating a positive feedback loop between cooperative preventive behavior and epidemic containment. Unfortunately, even with the rapid advancement of genetic sequencing technologies, the financial burden of achieving the sequencing rate necessary for our proposed analysis is quite high, and we cannot expect that we will have the same coverage in every pandemic. Deciding how much sequencing is actually needed for epidemic surveillance is currently an active research topic, as the cost-benefit trade-offs are still being debated [52]. Our analysis adds to this discussion by bringing a new potential benefit of dense genetic sequencing.
@@ Methods
and $\mathrm{ECS}_{c}$ is defined as the median of the $\mathrm{ECS}_{c,i}(t)$ values for all superspreading events $(c,i,t)$.
In Figure 4 and Supplementary Figure B.3, $\mathrm{ECS}_{c}$ values are compared with various exogenous variables (sequencing rate, attack rate, containment health index). These exogenous variables are computed for each country on a weekly basis based on publicly available datasets on the case counts [53], population counts [54], and the Oxford Containment Health Index [40]. Then, each superspreading event in the dataset is matched with the exogenous variables based on the time and country information. Finally, the plotted values are computed as the median of the exogenous variables of the superspreading events corresponding to index $c$ (which are also used to compute $\mathrm{ECS}_{c}$). See Supplementary Material A for a detailed explanation of these methodological choices in this section.
@@
Geometric Inhomogenous Random Graphs (GIRGs) were generated by sampling the spatial coordinates and the expected degrees of the nodes, and then connecting them by edges with a probability given by a kernel function, which is inversely proportional with the spatial distance, and assures the desired node degrees [55]. We used the Python implementation [56] for the sampling procedure with degree exponent $\tau=3.5$ and parameters $\alpha=2.3$, $C_{1}=0.8$. We tuned $C_{2}$ numerically to achieve the desired average degree (by default 3). Configuration models are generated by degree-preserving edge shuffling of the edges of the generated GIRG networks. SBMs were generated with blocks of size 50. The connection probabilities inside and between of the blocks were tuned so that for each node, half of it’s average degree were inside the block, and half of it’s average degree were outside the block. All synthetic networks had $10^{4}$ nodes, and we took the largest connected component if the network was not connected. We include a visualization of the size, degree distribution and average clustering coefficient of the generated networks in Supplementary Figure B.8.
@@
Data and Code Availability
All genome sequences and associated metadata are published in GISAID’s EpiCoV database. To view the contributors of each individual sequence with details such as accession number, Virus name, Collection date, Originating Lab and Submitting Lab and the list of Authors, visit 10.55876/gis8.240404rn. The MASZK survey data is available upon request from the authors. The code for the genetic data generation and analysis pipeline shown in Figure 1 (b) will be made available upon publication.
@@ Acknowledgements
We thank Eszter Ari, Andreas Bergthaler, and Tamás Stirling for their insightful comments and remarks. We also thank Júlia Koltai and Gergely Röst for their contribution to the MASZK survey data collection. GÓ was supported by the Swiss National Science Foundation, under grant number P500PT-211129. M.K. was supported by the CHIST-ERA project SAI: FWF I 5205-N; the SoBigData++ H2020-871042; SoBigData-PPP HORIZON-INFRA-2021-DEV-02 program under grant agreement No 101079043, and the National Laboratory for Health Security, Alfréd Rényi Institute, RRF-2.3.1-21-2022-00006.
@@ Appendix A Detailed explanation of superspreading event detection and ECS assignment
We present the distribution of $\mathrm{NormChange}_{c,i}(t+1)$ values for superspreading events, the $\mathrm{Baseline}_{c,i}(t+1)$ values for baseline events, along with the resulting ECS values in Belgium, in Figure A.1 (d). Since all country-variant pairs $c$ in our dataset had similarly broad, but unimodal ECS distributions as Figure A.1 (d), we focused on their median values denoted by $\mathrm{ECS}_{c}$ (the output of the pipeline in Figure 1 (b) of the main text). As the non-synonymous mutation rate of SARS-CoV-2 (we estimated a non-synonymous mutation probability of around 0.32 during a transmission event) was higher than the effective reproduction rate (often below 1.5), and collision clusters can be thought of as sub-critical spreading processes (with expected offspring number smaller than $(1-0.32)\cdot 1.5\approx 1$), it is no surprise that the median values of the $\mathrm{NormChange}_{c,i}(t+1)$ values are negative for both the superspreading events and the baselines. However, the sign of $\mathrm{ECS}_{c}$ adds non-trivial information. A positive $\mathrm{ECS}_{c}$ means that the normalized change of the number of genetic sequences in superspreading events was smaller than in the baseline, which suggests that the superspreading events led to fewer secondary infections than a similarly sized non-superspreading clusters of infectious individuals, i.e. the superspreading events were well-contained. Similarly, a negative $\mathrm{ECS}_{c}$ would suggest superspreading events that were not contained as well as the baselines in the same country during the same variant.
@@ B.1 Threshold for superspreading event Detection
To further strengthen the validity of our results, we present a robustness analysis on the threshold parameter. First, in Figure B.1 we show the number of detected superspreading events in various European countries as a function of the threshold parameter, if at least 15 superspreading events were detected (and therefore qualified for our analysis). As expected, the number of superspreading events is a monotone decreasing function of the threshold. Moreover, due to the log-scale it appears that the number of detected superspreading events decrease exponentially with the threshold, indicating that it is sufficient to perform the robustness analysis in a relatively narrow parameter range. We selected the interval [7,11] because a threshold of 11 only detects a minimal number of superspreading events in many countries, making them ineligible for our analysis, while a threshold of 7 results in a high number of superspreading events, potentially leading to an excessive number of false positives.
In Figures B.2-B.4, we recreated Figure 4 of the main text for each integer superspreading event detection threshold in the range [8,10], for all of the major SARS-CoV-2 variants. While there is some variability in the results for different thresholds (mostly due to new countries entering the dataset as the threshold decreases), besides the correlation between the $\mathrm{ECS}_{c}$ and the sampling date in the Delta wave also mentioned in the main text, the most significant correlations remain between the $\mathrm{ECS}_{c}$ and the Containment Health Index (CHI) in the Delta and the Omicron waves. These additional results further strengthen the conclusion made in the main text, that $\mathrm{ECS}_{c}$ is most correlated with the CHI (the most direct measure of human behavior) among the available exogenous variables, which includes potential confounding factors (sequencing rate, attack rate).
@@ B.3 Simulation parameters
Figure B.2: The figure shows how Figure 4 of the main text would look like if a threshold of 8 was chosen instead of the default value (9) for the Delta and Omicron waves. Grey background signifies a statistically significant correlation between the $\mathrm{ECS}_{c}$ and the exogenous variable.
Figure B.3: The figure shows how Figure 4 of the main text for the Delta and Omicron waves with the default threshold (9). Grey background signifies a statistically significant correlation between the $\mathrm{ECS}_{c}$ and the exogenous variable.
Figure B.4: The figure shows how Figure 4 of the main text would look like if a threshold of 10 was chosen instead of the default value (9) for the Delta and Omicron waves. Grey background signifies a statistically significant correlation between the $\mathrm{ECS}_{c}$ and the exogenous variable.
Figure B.5: The figure shows how Figure 4 of the main text would look like if $m=2$ chosen instead of the default value ($m=5$) when matching baseline events to superspreading events for the Delta and Omicron waves. Grey background signifies a statistically significant correlation between the $\mathrm{ECS}_{c}$ and the exogenous variable.
Figure B.6: The figure shows how Figure 4 of the main text would look like if $m=10$ chosen instead of the default value ($m=5$) when matching baseline events to superspreading events for the Delta and Omicron waves. Grey background signifies a statistically significant correlation between the $\mathrm{ECS}_{c}$ and the exogenous variable.
Figure B.7: ECS values computed on synthetically generated genetic sequence data similarly to Figure 3 of the main text, except with linear local and global awareness functions (see Methods for the precise function definition).

### v3

@@ Discussion
As with any trade-off, our proposed analysis comes with a number of limitations. We identify superspreading events based on simple thresholding of sequence counts, which is less accurate than manual contact tracing, where more metadata and more context about infection events can be taken into account. Consequently, we only compute highly aggregated statistics on the detected events. One ECS gives only very noisy information about the outcome of each superspreading event, and only the median of all ECSs, the MECS value has the statistical power to say anything about local awareness in region $c$. Since the amount of genetic sequences we have available since COVID-19 is unprecedented, and the new tools to analyze it are just being developed [23], our results too have to be confirmed by further research.
@@
From an operational perspective, by studying MECS values, we are able to measure how effectively different countries managed to contain superspreading events in different waves. We observe that this effectiveness is highly correlated with the containment policies implemented in each country, suggesting that stricter government policies could motivate the public to undertake stricter voluntary prevention methods. We envision that similar analyses will be used to evaluate the effectiveness of the implemented policies in future pandemics, potentially generating a positive feedback loop between cooperative preventive behavior and epidemic containment. Unfortunately, even with the rapid advancement of genetic sequencing technologies, the financial burden of achieving the sequencing rate necessary for our proposed analysis is quite high, and we cannot expect that we will have the same coverage in every pandemic. Deciding how much sequencing is actually needed for epidemic surveillance is currently an active research topic, as the cost-benefit trade-offs are still being debated [54]. Our analysis adds to this discussion by bringing a new potential benefit of dense genetic sequencing.
@@ Methods
and MECS for country $c$ is defined as the median of the $\mathrm{ECS}_{c,i}(t)$ values for all superspreading events $(c,i,t)$ in $c$.
In Figure 4 and Supplementary Figure B.3, MECS values are compared with various exogenous variables (sequencing rate, attack rate, Containment Health Index). These exogenous variables are computed for each country on a weekly basis based on publicly available datasets on the case counts [55] and the Oxford Containment Health Index [41]. Then, each superspreading event in the dataset is matched with the exogenous variables based on the time and country information. Finally, the plotted values are computed as the median of the exogenous variables of the superspreading events corresponding to index $c$ (which are also used to compute MECS). See Supplementary Material A for a detailed explanation of these methodological choices in this section.
@@
Geometric Inhomogenous Random Graphs (GIRGs) were generated by sampling the spatial coordinates and the expected degrees of the nodes, and then connecting them by edges with a probability given by a kernel function, which is inversely proportional with the spatial distance, and assures the desired node degrees [56]. To sample networks with a heterogeneous degree distribution and geometric properties [57], we set the degree exponent to $\tau=3.5$ and the parameters to $\alpha=2.3$, $C_{1}=0.8$. We tuned $C_{2}$ numerically to achieve the desired average degree (by default 3). Configuration models are generated by degree-preserving edge shuffling of the edges of the generated GIRG networks. SBMs were generated with blocks of size 50. The connection probabilities inside and between of the blocks were tuned so that for each node, half of it’s average degree were inside the block, and half of it’s average degree were outside the block. All synthetic networks had $10^{4}$ nodes, and we took the largest connected component if the network was not connected. We include a visualization of the size, degree distribution and average clustering coefficient of the generated networks in Supplementary Figure B.8.
@@
For each set of parameters, we simulated $n=200$ epidemic processes with different random seeds. When the underlying network was synthetic, it was generated with the same random seed as the epidemic process prior to the simulation.
@@
Data Availability
The MASZK survey data used to generate Figure 1 is shared in Supplementary Data 1. The simulated data in Figure 3 is available in Supplementary Data 2. All genome sequences and associated metadata are published in GISAID’s EpiCoV database. To view the contributors of each individual sequence with details such as accession number, Virus name, Collection date, Originating Lab and Submitting Lab and the list of Authors, visit 10.55876/gis8.240404rn. For the reported COVID-19 case numbers, we used the “JHU CSSE COVID-19 Data” available at https://github.com/CSSEGISandData/COVID-19 applying a 7-day rolling average and outlier detection to ensure data consistency and reliability. An intermediate dataset containing the accession numbers of the sequences, the computed ECS values, and various additional metadata of the collision clusters corresponding to the detected superspreading events is shared in Supplementary Data 3.
Code Availability
The code for the genetic data generation and analysis pipeline shown in Figure 1 (b) is available via Code Ocean at https://doi.org/10.24433/CO.1189503.v1.
@@ Acknowledgements
We thank Eszter Ari, Andreas Bergthaler, and Tamás Stirling for their insightful comments and remarks. We also thank Júlia Koltai and Gergely Röst for their contribution to the MASZK survey data collection. We gratefully acknowledge all data contributors, i.e., the Authors and their Originating laboratories responsible for obtaining the specimens, and their Submitting laboratories for generating the genetic sequence and metadata and sharing via the GISAID Initiative, on which this research is based. GÓ was primarily supported by the Swiss National Science Foundation, under grant number P500PT-211129, with additional funding from the Austrian Science Fund (FWF) Cluster of Excellence “Microbiomes drive planetary health” (10.55776/COE7). M.K. was supported by the CHIST-ERA project SAI: FWF I 5205-N; the SoBigData++ H2020-871042; SoBigData-PPP HORIZON-INFRA-2021-DEV-02 program under grant agreement No 101079043, and the National Laboratory for Health Security, Alfréd Rényi Institute, RRF-2.3.1-21-2022-00006.
@@
Competing interests
The authors declare no competing interests.
@@ Appendix A Detailed explanation of superspreading event detection and ECS assignment
We present the distribution of $\mathrm{NormChange}_{c,i}(t+1)$ values for superspreading events, the $\mathrm{Baseline}_{c,i}(t+1)$ values for baseline events, along with the resulting ECS values in Belgium, in Figure A.1 (d). Since all country-variant pairs $c$ in our dataset had similarly broad, but unimodal ECS distributions as Figure A.1 (d), we focused on their median values denoted by MECS. As the non-synonymous mutation rate of SARS-CoV-2 (we estimated a non-synonymous mutation probability of around 0.32 during a transmission event) was higher than the effective reproduction rate (often below 1.5), and collision clusters can be thought of as sub-critical spreading processes (with expected offspring number smaller than $(1-0.32)\cdot 1.5\approx 1$), it is no surprise that the median values of the $\mathrm{NormChange}_{c,i}(t+1)$ values are negative for both the superspreading events and the baselines. However, the sign of MECS adds non-trivial information. A positive MECS means that the normalized change of the number of genetic sequences in superspreading events was smaller than in the baseline, which suggests that the superspreading events led to fewer secondary infections than a similarly sized non-superspreading clusters of infectious individuals, i.e. the superspreading events were well-contained. Similarly, a negative MECS would suggest superspreading events that were not contained as well as the baselines in the same country during the same variant.
@@ B.1 Threshold for superspreading event Detection
To further strengthen the validity of our results, we present a robustness analysis on the threshold parameter. First, in Figure B.1 we show the number of detected superspreading events in various European countries as a function of the threshold parameter, if at least 15 superspreading events were detected (and therefore qualified for our analysis). As expected, the number of superspreading events is a monotone decreasing function of the threshold. Moreover, due to the log-scale it appears that the number of detected superspreading events decrease exponentially with the threshold, indicating that it is sufficient to perform the robustness analysis in a relatively narrow parameter range. We selected the interval [7,11] because a threshold of 11 only detects a minimal number of superspreading events in many countries, making them ineligible for our analysis, while a threshold of 7 results in a high number of superspreading events, potentially leading to an excessive number of false positives. For completeness we also report the number of superspreading events corresponding to each MECS value shown in Figure 4 of the main text in Table B.1.
(a) Delta variant (b) Omicron BA.1 variant (c) Omicron all variants
country number of ST country number of ST country number of ST
Sweden 79 United Kingdom 717 Sweden 48
Switzerland 35 Sweden 23 Ireland 68
Slovenia 25 Belgium 23 United Kingdom 1163
Denmark 529 Germany 66 Denmark 139
United Kingdom 826 Slovenia 42 Belgium 35
Norway 47 Germany 199
Belgium 17 Slovenia 51
Ireland 37
Germany 67
Table B.1: The number of superspreading events corresponding to each MECS value in Figure 4.
In Figures B.2-B.4, we recreated Figure 4 of the main text for each integer superspreading event detection threshold in the range [8,10], for all of the major SARS-CoV-2 variants. While there is some variability in the results for different thresholds (mostly due to new countries entering the dataset as the threshold decreases), besides the correlation between the MECS and the sampling date in the Delta wave also mentioned in the main text, the most significant correlations remain between the MECS and the Containment Health Index (CHI) in the Delta and the Omicron waves. These additional results further strengthen the conclusion made in the main text, that MECS is most correlated with the CHI (the most direct measure of human behavior) among the available exogenous variables, which includes potential confounding factors (sequencing rate, attack rate).
@@ B.3 Simulation parameters
Figure B.2: The figure shows how Figure 4 of the main text would look like if a threshold of 8 was chosen instead of the default value (9) for the Delta and Omicron waves. Grey background signifies a statistically significant correlation between the MECS and the exogenous variable. Error bar definitions and all plotting parameters are the same as in Figure 4.
Figure B.3: The figure shows how Figure 4 of the main text for the Delta and Omicron waves with the default threshold (9). Grey background signifies a statistically significant correlation between the MECS and the exogenous variable. Error bar definitions and all plotting parameters are the same as in Figure 4.
Figure B.4: The figure shows how Figure 4 of the main text would look like if a threshold of 10 was chosen instead of the default value (9) for the Delta and Omicron waves. Grey background signifies a statistically significant correlation between the MECS and the exogenous variable. Error bar definitions and all plotting parameters are the same as in Figure 4.
Figure B.5: The figure shows how Figure 4 of the main text would look like if $m=2$ chosen instead of the default value ($m=5$) when matching baseline events to superspreading events for the Delta and Omicron waves. Grey background signifies a statistically significant correlation between the MECS and the exogenous variable. Error bar definitions and all plotting parameters are the same as in Figure 4.
Figure B.6: The figure shows how Figure 4 of the main text would look like if $m=10$ chosen instead of the default value ($m=5$) when matching baseline events to superspreading events for the Delta and Omicron waves. Grey background signifies a statistically significant correlation between the MECS and the exogenous variable. Error bar definitions and all plotting parameters are the same as in Figure 4.
Figure B.7: ECS values computed on synthetically generated genetic sequence data similarly to Figure 3 of the main text, except with linear local and global awareness functions (see Methods for the precise function definition). The sample size, error bar definitions and all plotting parameters are the same as in Figure 3.